### v1

# OneOcc: Semantic Occupancy Prediction for Legged Robots with a Single Panoramic Camera

Abstract
Robust 3D semantic occupancy is crucial for legged/humanoid robots, yet most semantic scene completion (SSC) systems target wheeled platforms with forward-facing sensors. We present OneOcc, a vision-only panoramic SSC framework designed for gait-introduced body jitter and $360^{\circ}$ continuity. OneOcc combines: (i) Dual-Projection fusion (DP-ER) to exploit the annular panorama and its equirectangular unfolding, preserving $360^{\circ}$ continuity and grid alignment; (ii) Bi-Grid Voxelization (BGV) to reason in Cartesian and cylindrical-polar spaces, reducing discretization bias and sharpening free/occupied boundaries; (iii) a lightweight decoder with Hierarchical AMoE-3D for dynamic multi-scale fusion and better long-range/occlusion reasoning; and (iv) plug-and-play Gait Displacement Compensation (GDC) learning feature-level motion correction without extra sensors. We also release two panoramic occupancy benchmarks: QuadOcc (real quadruped, first-person $360^{\circ}$) and Human360Occ (H3O) (CARLA human-ego $360^{\circ}$ with RGB, Depth, semantic occupancy; standardized within-/cross-city splits). OneOcc sets new state-of-the-art (SOTA): on QuadOcc it beats strong vision baselines and popular LiDAR ones; on H3O it gains +3.83 mIoU (within-city) and +8.08 (cross-city). Modules are lightweight, enabling deployable full-surround perception for legged/humanoid robots. Datasets and code will be publicly available at https://github.com/MasterHow/OneOcc.
1 Introduction
Figure 1: Teaser. (a) Single-sensor panoramic semantic scene completion (SSC) on a quadruped: an omnidirectional camera observes full $360^{\circ}$ with gait-induced motion; OneOcc outputs voxelized semantic occupancy. (b) On QuadOcc, OneOcc reaches 20.56 mIoU, surpassing the best LiDAR baseline (LMSCNet, $18.44$) and vision baseline (MonoScene, $19.19$). On H3O, OneOcc attains 37.29/32.23 mIoU under within-/cross-city, outperforming the best vision baselines ($33.46/24.15$).
A holistic 3D understanding of nearby surroundings is fundamental for robotic autonomy in navigation [wang2024omega, hoeller2022neural, wang2024he_nav, wang2024agrnav, patel2025hierarchical, li2024stereonavnet], locomotion [niijima2025realtime_multi_plane, dong2025marg, escontrela2025gaussgym, chen2025learning_traversal], manipulation [kartmann2021semantic, dengler2025efficient_manipulation, li2025manipdreamer3d, wang2025odyssey], and human-robot interaction [wu2025embodiedocc, wang2025embodiedocc++, liu2024revisit_human_scene_interaction]. Among 3D perception tasks, Semantic Scene Completion (SSC)—predicting complete volumetric geometry and semantics from partial observations—has become central [lai2024rtonet, cao2024slcf_sequential, yan2024pointssc, zhang2025roboocc], with most progress driven by autonomous driving where forward-facing sensors on stable wheeled platforms yield benign motion and sensing.
Data Domain
Datasets Camera Sur. Plat. Move. Env. Weath. Lighting Modality Volume Size Scenes Frames Anno.
NYUv2 [nyu] Pinhole [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Indoor [Uncaptioned image] [Uncaptioned image] C&D $(144{\times}240{\times}240)$ 1.4K 1.4K [Uncaptioned image]
ScanNet [scannet] Pinhole [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Indoor [Uncaptioned image] [Uncaptioned image] C&D $(31{\times}62{\times}62)$ 1.5K 1.5K [Uncaptioned image]
SUNCG [song2017semantic] Pinhole [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Indoor [Uncaptioned image] [Uncaptioned image] C&D $(144{\times}240{\times}240)$ 46K 140K [Uncaptioned image]
SemanticPOSS [semanticposs] n.a. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Outdoor [Uncaptioned image] [Uncaptioned image] L $(32{\times}256{\times}256)$ 6 3K [Uncaptioned image]
SemanticKITTI [semantickitti] Pinhole [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Outdoor [Uncaptioned image] [Uncaptioned image] C&L $(32{\times}256{\times}256)$ 22 9K [Uncaptioned image]
nuScenes [wang2023openoccupancy] Pinhole+Fisheye [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Outdoor [Uncaptioned image] [Uncaptioned image][Uncaptioned image][Uncaptioned image] C&L $(40{\times}512{\times}512)$ 850 200K [Uncaptioned image]
Occ3D-Waymo [tian2024occ3d] Pinhole+Fisheye [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Outdoor [Uncaptioned image] [Uncaptioned image][Uncaptioned image][Uncaptioned image] C&L $(32{\times}200{\times}200)$ 1K 200K [Uncaptioned image]
SSCBench-KITTI-360 [li2024sscbench] Fisheye [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Outdoor [Uncaptioned image] [Uncaptioned image] C&L $(32{\times}200{\times}200)$ 11 91K [Uncaptioned image]
ScribbleSC [wang2024label_efficient] Pinhole [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Outdoor [Uncaptioned image] [Uncaptioned image] C&L $(32{\times}256{\times}256)$ 11 9K [Uncaptioned image]
OmniHD-Scenes [zheng2024omnihd] Pinhole [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Outdoor [Uncaptioned image] [Uncaptioned image][Uncaptioned image] C&L&R 200 60K [Uncaptioned image][Uncaptioned image]
WildOcc [zhai2024wildocc] Pinhole [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Outdoor [Uncaptioned image] [Uncaptioned image] C&L $(40{\times}100{\times}100)$ 5 10K [Uncaptioned image]
DSEC-SSC [guo2025event] Pinhole [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Outdoor [Uncaptioned image] [Uncaptioned image][Uncaptioned image] C&L&E $(16{\times}128{\times}128)$ 6 3K [Uncaptioned image]
ORAD-3D [min2025advancing_orad] Pinhole [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Outdoor [Uncaptioned image] [Uncaptioned image][Uncaptioned image][Uncaptioned image] C&L 145 58K [Uncaptioned image]
Co3SOP [wu2025synthetic_v2x] Pinhole [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Outdoor [Uncaptioned image] [Uncaptioned image] C&L $(70{\times}1000{\times}1000)$ [Uncaptioned image]
Human360Occ (ours) Panoramic [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Outdoor [Uncaptioned image] [Uncaptioned image][Uncaptioned image][Uncaptioned image] C&D $(16{\times}128{\times}128)$ 160 8K [Uncaptioned image]
QuadOcc (ours) Panoramic [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Outdoor [Uncaptioned image] [Uncaptioned image][Uncaptioned image][Uncaptioned image] C&L $(8{\times}64{\times}64)$ 10 24K [Uncaptioned image][Uncaptioned image]
Table 1: Typical datasets for Semantic Scene Completion (SSC). Abbreviations: [Uncaptioned image] (Autonomous Car), [Uncaptioned image] (Quadruped Robot), [Uncaptioned image] (Vision Foundation Model), [Uncaptioned image] (Human), [Uncaptioned image] (Synthetic), [Uncaptioned image] (Internet images/videos), [Uncaptioned image] (Wheels), [Uncaptioned image] (Gait), [Uncaptioned image] (Stationary), [Uncaptioned image] (Day), [Uncaptioned image] (Dusk), [Uncaptioned image] (Night), Sur. (Surround), Plat. (Platform), Move. (Movement), Env. (Environment), Weath. (Weather diversity), Anno. (Annotation), C (Camera), D (Depth), L (LiDAR), R (Radar).
Why legged and humanoid platforms are different. Directly transferring SSC is non-trivial: agile gaits cause body jitter that corrupts evidence and breaks temporal coherence; these platforms require full $360^{\circ}$ situational awareness in cluttered, uneven terrain; and tight payload/power budgets favor lightweight, single-sensor, low-latency solutions. Hence, we need a panoramic SSC pipeline explicitly robust to gait-induced motion.
Panoramic SSC with a single sensor. A single panoramic camera provides compact, true $360^{\circ}$ coverage but introduces annular distortions and seam artifacts after unwrapping [lin2025one_flight]—issues poorly handled by perspective-camera SSC pipelines. This motivates a design that preserves raw panoramic cues while exploiting grid-friendly projections.
Method: OneOcc. We propose OneOcc, a vision-only panoramic SSC framework for severe body jitter and $360^{\circ}$ continuity, integrating: (i) Dual-Projection fusion (DP-ER) to jointly process the raw annular panorama and its equirectangular unfolding; (ii) Bi-Grid Voxelization (BGV) that reasons in Cartesian and polar/cylindrical voxel spaces to reduce discretization bias and better match panoramic geometry; (iii) a lightweight decoder with Hierarchical AMoE-3D for dynamic multi-scale 3D fusion; and (iv) a plug-and-play Gait Displacement Compensation (GDC) that learns feature-level jitter correction without extra sensors.
Figure 2: OneOcc pipeline. Calibrated unwrapping maps the raw panorama to equirectangular. Two 2D encoders (DP-ER) feed GDC for feature-level gait compensation (optional on H3O). At three scales, BGV builds Cartesian and polar/cylindrical voxel grids; voxel-centroid features are sampled to each view (Ca/Po $\rightarrow$ Equi/Raw) and fused. A lightweight 3D decoder with Hierarchical AMoE-3D aggregates multi-scale context, followed by a head predicting full-surround occupancy. Abbrev.: Ca/Po = Cartesian/Polar; Equi = equirectangular.
Benchmarks: QuadOcc and Human360Occ (H3O). To advance panoramic SSC on legged/humanoid robots, we release two complementary benchmarks. QuadOcc is a real first-person $360^{\circ}$ dataset on a quadruped within a campus domain, with standardized day/dusk/night coverage and semi-automatic ground truth (multi-frame LiDAR aggregation, Grounded-SAM [ren2024grounded] initialization, targeted manual fixes). It contains 10 scenes and 24K frames, uses 6 semantic categories for training. The occupancy uses 0.4 m voxels on a $64\!\times\!64\!\times\!8$ grid—reflecting embodied compute and lower speed than cars. Human360Occ (H3O) is CARLA-based human-ego $360^{\circ}$ with simulated gait, 160 sequences / 8K frames over 16 maps and multiple weathers/times; each frame provides RGB, metric depth, semantic occupancy at $\times64$$\times$8 and $\times128$$\times$16, and pose. It supports within-city (per-map 8:28{:}2 train/val) and cross-city (default 12/412{/}4 maps) splits.
Results in brief. On QuadOcc, OneOcc reaches 20.56 mIoU, surpassing the best LiDAR baseline (LMSCNet, $18.44$; +2.12+2.12 mIoU, $11.5\%$) and the best vision baseline (MonoScene, $19.19$; +1.37+1.37 mIoU). On H3O, it attains 37.29/32.23 mIoU under within-/cross-city, improving over the best vision baselines ($33.46/24.15$) by +3.83+3.83 and +8.08+8.08 mIoU (up to $33.5\%$ relative); see Fig. 1(b).
Contributions.
  • Method. OneOcc integrates DP-ER, BGV, Hierarchical AMoE-3D, and GDC to deliver jitter-robust, full-surround SSC with lightweight computation.
  • Datasets. Two panoramic occupancy benchmarks for legged robots: QuadOcc (real quadruped, $360^{\circ}$, $10$ scenes/2424K frames, $6$ categories, 0.40.4m voxels on $64\!\times\!64\!\times\!8$) and H3O (CARLA human-ego $360^{\circ}$, $160$ seq/88K frames on $16$ maps, per-frame RGB/Depth/voxels/pose, two voxel resolutions).
  • Evaluation. Extensive benchmarking shows vision-only OneOcc rivals or surpasses classic LiDAR-based methods under realistic legged/humanoid settings.
2 Related Work
[…]
Semantic scene understanding for legged robots. Legged autonomy motivates perception tailored to agile motion and strict payload budgets. Prior efforts include vision-based terrain reconstruction [zhang2022vision], large-scale datasets [patel2025tartanground], and traversability learning [deng2021vision, aegidius2025watch, oh2024trip]; complementary modules cover robust tracking [cao2025siamese, xin2024robotic] and instruction grounding [mei2024quadrupedgpt, han2025space]. With increasing evidence that limited FoV hinders agility [li2024move, cheng2024quadruped, wang2025omni], panoramic sensors are adopted for omnidirectional tracking [luo2025omnidirectional], multimodal odometry [li2025limo], and even satellite communication on the move [liu2025starlink]. Some works study probabilistic semantic mapping [chen2025particle], and humanoid systems explore panoramic/occupancy perception [cui2025humanoid, zhang2025humanoidpano, zhang2025roboocc, zhang2025occupancy_world_models_robots]. However, dense 3D semantic occupancy from a single panoramic camera on dynamic legged platforms is still largely open. We address this by releasing QuadOcc and Human360Occ and proposing OneOcc, a lightweight panoramic occupancy framework built around dual-projection fusion, bi-grid voxelization, hierarchical 3D MoE fusion, and gait displacement compensation.
Figure 3: GDC and AMoE-3D Fusion Module. (a) GDC regresses $\Delta=(dx,dy)$ from 2D features via GAP+Linear (zero–init), applies scale/shift to multi–scale view coordinates $\{C_{s}\}$ for voxel–centroid sampling, and yields $\{F_{3d}\}_{V2V}$. (b) AMoE-3D fuses Po2Ca-sampled polar features with Cartesian ones via channel alignment, concatenation, and reduction, applies 3D channel- and spatial attention, and performs MoE–Fuse3D with $K$ experts and GradEnergy3D–Softmax gating: $y=\sum_{k}P_{k}\odot E_{k}(x)$. The fused $F^{\mathrm{Fusion}}_{3d}$ feeds the 3D decoder.
3 Method
3.1 Problem Formulation
Given a panoramic RGB image $\mathbf{I}\!\in\!\mathbb{R}^{H\times W\times 3}$ and calibrated intrinsics/extrinsics, we aim to predict a dense 3D semantic occupancy $\mathbf{S}\!\in\!\{0,\ldots,C\}^{X\times Y\times Z}$. We produce per-voxel logits $\mathbf{Z}\!\in\!\mathbb{R}^{X\times Y\times Z\times C}$, and the final label is $\hat{s}_{x,y,z}=\arg\max_{c}\mathbf{Z}_{x,y,z,c}$.
3.2 Overview
[…]
3.3 Unwrapping and Dual-Projection Encoders
Motivation.
Equirectangular views preserve azimuthal continuity that is convolution-friendly; the raw panoramic annulus retains native geometry and fine textures. Running both in parallel compensates the equator–pole trade-off.
Method.
We calibrate the omnidirectional panoramic annular lens (PAL) camera with the Taylor polynomial model [scaramuzza2006toolbox]. The unwrapping samples the raw annulus at coordinates given by the calibrated inverse mapping:
$\mathbf{I}^{\mathrm{equi}}(u,v)\;=\;\mathbf{I}^{\mathrm{raw}}\!\left(\,\mathcal{U}^{-1}(u,v;\boldsymbol{\kappa})\,\right),$ (1)
where $\boldsymbol{\kappa}=\{a_{i},\,u_{0},\,v_{0},\,\mathbf{A}\}$ are the Taylor coefficients and principal point plus a $2{\times}2$ affine matrix (scale/skew), and $(u,v)\!\in\![0,W)\!\times\![0,H)$. For the equirectangular ray with spherical angles $(\phi,\theta)$:
$\phi\;$ $=\;\tfrac{2\pi}{W}\,u-\pi,\qquad\theta\;=\;\tfrac{\pi}{2}-\tfrac{\pi}{H}\,v,$ (2)
$r(\theta)\;$ $=\;\sum_{i=0}^{N}a_{i}\,\theta^{\,i},\qquad\begin{bmatrix}u_{\text{raw}}\\[2.0pt] v_{\text{raw}}\end{bmatrix}=\begin{bmatrix}u_{0}\\[2.0pt] v_{0}\end{bmatrix}+\mathbf{A}\,r(\theta)\begin{bmatrix}\cos\phi\\[2.0pt] \sin\phi\end{bmatrix}$. (3)
Two encoders then produce multi-scale features:
$\{\mathbf{F}^{\mathrm{equi}}_{1/s}\}_{s\in\{1,4,8,16\}},\qquad\{\mathbf{F}^{\mathrm{raw}}_{1/s}\}_{s\in\{1,4,8,16\}},$ (4)
with GroupNorm and $1{\times}1$ bottleneck alignment to standardize channels.
Why it helps.
For legged robots, the equatorial ring dominates motion priors while polar areas are highly distorted. Using the Taylor unwrapping respects PAL optics, preserving ring continuity in ER space and local textures in the raw annulus, providing more stable cues for lifting.
Table 2: Semantic scene completion results on the QuadOcc validation set. All models are (re)trained by us on QuadOcc. Bold = best, underline = second best (within the LiDAR-based block; Vision-based block marked separately.)
Method Input
$\blacksquare$ vehicle (0.90%)
$\blacksquare$ pedestrian (2.09%)
$\blacksquare$ road (52.34%)
$\blacksquare$ building (14.00%)
$\blacksquare$ vegetation (22.82%)
$\blacksquare$ terrain (7.85%)
Precision $\uparrow$ Recall $\uparrow$ IoU $\uparrow$ mIoU $\uparrow$
LiDAR-based
SSCNet [song2017semantic] L 0.00 0.04 44.34 16.06 22.41 4.77 63.42 65.82 47.71 14.60
SSCNet-full [song2017semantic] L 0.71 0.04 55.14 14.93 19.66 7.76 78.81 60.03 51.69 16.38
LMSCNet [roldao2020lmscnet] L 0.88 0.20 57.02 16.45 24.39 11.70 79.70 64.49 55.40 18.44
OccRWKV [wang2024occrwkv] L 2.05 0.35 54.17 10.29 24.92 8.38 71.98 62.80 50.47 16.69
Vision-based
VoxFormer-S [li2023voxformer] P+D 0.31 1.67 36.83 6.48 7.09 3.81 45.96 69.95 38.38 9.36
SGN [mei2023camera] P+D 7.39 1.88 53.35 9.06 21.95 9.39 59.23 67.41 46.05 17.17
VoxFormer-S [li2023voxformer] P+L 0.24 1.66 38.51 10.90 11.88 4.45 51.29 75.84 44.08 11.27
VoxFormer-T [li2023voxformer] P+L 0.30 1.65 40.21 14.48 15.86 4.64 56.94 77.10 48.70 12.86
SGN [mei2023camera] P+L 11.62 2.47 53.06 15.60 25.67 9.91 73.70 59.62 49.16 19.72
OccFormer [zhang2023occformer] P 0.29 0.37 49.46 10.36 15.00 2.64 45.79 76.99 40.28 13.02
MonoScene [cao2022monoscene] P 8.15 1.59 55.66 12.88 26.10 10.78 62.10 69.28 48.69 19.19
OneOcc (Ours) P 12.16 2.86 54.41 16.03 24.91 13.01 66.69 64.74 48.92 20.56
  • $\dagger$
    Uses a different sensor modality from the original paper (here: adding LiDAR).
  • Input legend: L = LiDAR; P = Panoramic camera; D = monocular Depth (pred.).
Table 3: Semantic scene completion using Panoramas on H3O (Human360Occ). Two vertically-stacked panels report results on HOMO (within-city) and HETER (cross-city) splits. Header frequencies are computed on non-empty voxels. Input legend: P = Panoramic camera; Dpred. = predicted monocular depth; Dgt = ground-truth depth.
HOMO (within-city)
Method Input
$\blacksquare$ road (34.04%)
$\blacksquare$ sidewalk (26.24%)
$\blacksquare$ building (10.00%)
$\blacksquare$ vegetation (24.02%)
$\blacksquare$ car (2.92%)
$\blacksquare$ truck (0.99%)
$\blacksquare$ bus (0.25%)
$\blacksquare$ two_wheeler (0.07%)
$\blacksquare$ person (0.97%)
$\blacksquare$ pole (0.50%)
Precision Recall IoU mIoU
VoxFormer-[li2023voxformer] P+Dpred. 22.27 17.44 2.86 5.65 0.87 0.14 0.00 0.00 61.46 0.18 40.21 59.49 31.57 11.09
VoxFormer-[li2023voxformer] P+Dgt 35.84 25.86 12.32 9.90 3.73 0.70 0.19 0.17 64.79 0.93 59.91 71.12 48.19 15.44
SGN-[mei2023camera] P+Dpred. 54.55 44.61 15.51 35.16 23.91 7.68 4.52 15.77 65.24 19.46 61.82 64.36 46.05 28.64
SGN-[mei2023camera] P+Dpred. 54.81 47.44 15.64 33.22 24.23 8.53 2.62 13.51 65.88 19.72 62.28 63.72 45.98 28.56
OccFormer [zhang2023occformer] P 54.11 46.41 15.14 32.84 17.57 4.30 0.00 1.65 65.08 11.72 57.30 69.94 45.98 24.88
MonoScene [cao2022monoscene] P 62.45 57.18 19.97 38.91 29.76 13.77 5.03 16.76 70.15 20.67 68.78 71.66 54.08 33.46
OneOcc (Ours) P 63.82 62.92 22.86 41.03 33.74 17.79 6.47 26.58 71.13 26.55 69.14 74.18 55.73 37.29
HETER (cross-city)
Method Input
$\blacksquare$ road (34.04%)
$\blacksquare$ sidewalk (26.24%)
$\blacksquare$ building (10.00%)
$\blacksquare$ vegetation (24.02%)
$\blacksquare$ car (2.92%)
$\blacksquare$ truck (0.99%)
$\blacksquare$ bus (0.25%)
$\blacksquare$ two_wheeler (0.07%)
$\blacksquare$ person (0.97%)
$\blacksquare$ pole (0.50%)
Precision Recall IoU mIoU
VoxFormer-[li2023voxformer] P+Dpred. 19.02 15.82 2.84 5.86 0.95 0.10 0.00 0.01 61.62 0.10 36.68 52.00 27.40 10.63
VoxFormer-[li2023voxformer] P+Dgt 32.55 25.95 8.98 9.93 4.68 0.44 0.00 0.00 65.45 0.26 59.08 61.21 42.99 14.82
SGN-[mei2023camera] P+Dpred. 42.57 23.94 9.42 15.84 18.22 10.21 1.55 3.11 63.44 11.93 45.84 47.11 30.26 20.02
SGN-[mei2023camera] P+Dpred. 43.47 24.70 9.19 20.27 17.21 8.21 1.85 5.08 61.10 14.89 47.82 50.98 32.76 20.60
OccFormer [zhang2023occformer] P 47.00 39.01 10.14 24.85 11.84 1.52 0.05 0.77 66.60 6.93 53.12 63.57 40.73 20.87
MonoScene [cao2022monoscene] P 49.68 41.41 11.04 20.57 20.34 10.36 1.98 2.79 69.68 13.69 67.39 55.00 43.44 24.15
OneOcc (Ours) P 58.60 48.28 16.34 25.98 30.68 19.35 12.87 14.16 72.89 23.11 67.89 64.77 49.58 32.23
3.4 Bi-Grid Voxelization and View2View Sampling
Motivation.
Cartesian voxels are accurate for near-field contact geometry (footholds, obstacles), whereas polar/cylindrical voxels preserve azimuthal continuity that matches panoramic imaging. In a $360^{\circ}$ camera, the equirectangular horizontal axis is linearly related to azimuth $\phi$; cylindrical voxels $(r,\varphi,z)$ therefore form a natural, camera-aligned parameterization where $\varphi\!\approx\!\phi$ and $r$ correlates with depth, yielding uniform sampling along the ring and reduced far-field aliasing. Combining Cartesian and cylindrical grids reduces quantization bias and balances near/far evidence under legged motion.
Method.
We define voxel centroids
$\left\{\begin{aligned} \mathbf{c}^{\mathrm{Ca}}_{ijk}\;&=\;(x_{i},y_{j},z_{k}),\\ \mathbf{c}^{\mathrm{Po}}_{pqk}\;&=\;\big(r_{p}\cos\varphi_{q},\;r_{p}\sin\varphi_{q},\;z_{k}\big).\end{aligned}\right$. (5)
For view $v\!\in\!\{\mathrm{equi},\mathrm{raw}\}$ and scale $s$, we lift features by bilinear sampling at the projected pixel $\pi_{v}(\cdot)$:
$\left\{\begin{aligned} \mathbf{V}^{\mathrm{Ca}}_{s}(\mathbf{c})\;&=\;\mathrm{bilinear}\!\left(\mathbf{F}^{v}_{1/s},\,\pi_{v}(\mathbf{c};\boldsymbol{\kappa})\right),\\ \mathbf{V}^{\mathrm{Po}}_{s}(\mathbf{c})\;&=\;\mathrm{bilinear}\!\left(\mathbf{F}^{v}_{1/s},\,\pi_{v}(\mathbf{c};\boldsymbol{\kappa})\right).\end{aligned}\right$. (6)
We fuse scales with per-voxel convex weights:
$\left\{\begin{aligned} \mathbf{V}^{\mathrm{Ca}}\;&=\;\sum_{s}\alpha^{\mathrm{Ca}}_{s}\odot\mathbf{V}^{\mathrm{Ca}}_{s},\qquad\sum_{s}\alpha^{\mathrm{Ca}}_{s}=1,\\ \mathbf{V}^{\mathrm{Po}}\;&=\;\sum_{s}\alpha^{\mathrm{Po}}_{s}\odot\mathbf{V}^{\mathrm{Po}}_{s},\qquad\sum_{s}\alpha^{\mathrm{Po}}_{s}=1.\end{aligned}\right$. (7)
To inject polar context into the Cartesian stream, we pre-compute cross-grid indices $\{\mathcal{J}_{\ell}\}_{\ell\in\{1,2,4\}}$ and form
$\widetilde{\mathbf{V}}^{\mathrm{Ca}}_{\ell}\;=\;\mathrm{Align}_{1\times 1\times 1}\!\big(\,\mathbf{V}^{\mathrm{Po}}_{\ell}[\mathcal{J}_{\ell}]\,\big)\;\|\;\mathbf{V}^{\mathrm{Ca}}_{\ell}.$ (8)
Why it helps.
On legged robots, near-field foothold safety leverages Cartesian precision while far-field loop/context exploits cylindrical efficiency. The camera-aligned polar stream improves azimuthal coherence and reduces spherical-distortion artifacts; the fused bi-grid volume balances near/far evidence under gait jitters.
3.5 Gait Displacement Compensation
Motivation.
Gait shocks induce phase errors; correcting them after lifting acts on already voxel-quantized evidence. Compensating before lifting—at sampling coordinates—avoids quantization and is cheaper.
Method.
For each scale and projection, we regress a 2D displacement $\Delta_{s}\!=\!(d_{x},d_{y})$ with a zero-initialized head:
$\Delta_{s}\;=\;\mathrm{Linear}_{0}\!\Big(\mathrm{GAP}(\mathbf{F}^{v}_{1/s})\Big),$ (9)
where $\mathrm{Linear}_{0}$ has all weights and bias set to zero so the initial warp is identity (cf. zero-conv [zhang2023adding]). Before lifting we correct and resample:
$\widehat{\mathbf{p}}\;$ $=\;\pi_{v}(\mathbf{c};\boldsymbol{\kappa})\;+\;\Delta_{s},$ (10)
[…]
Here, $v\!\in\!\{\text{unwrapped},\text{spherical}\}$ indexes the projection path; $\pi_{v}(\mathbf{c};\boldsymbol{\kappa})$ projects a 3D voxel centroid $\mathbf{c}$ to image pixels using the calibrated Taylor camera with parameters $\boldsymbol{\kappa}$. $\widehat{\mathbf{p}}$ is the corrected sampling pixel; $\mathrm{bilinear}(\cdot)$ samples the 2D feature map $\mathbf{F}^{v}_{1/s}$, and $\mathbf{V}^{(\cdot)}_{s}$ denotes the lifted 3D feature.
Why it helps.
GDC routes the phase error back to 2D, preventing mis-attributed voxels. The zero-init head stabilizes early training and leaves features unperturbed when motion is tiny ($\Delta_{s}\!\approx\!0$).
3.6 Hierarchical AMoE-3D
Motivation.
Panoramic scenes are anisotropic: strong azimuthal variation vs. vertical structure, and large near/far scale disparity. We thus (i) inject polar cues hierarchically (coarse-to-fine), and (ii) apply voxel-wise expert selection to prevent over-smoothing in flat regions while enhancing edges/contacts critical for locomotion.
Method.
The 3D decoder is a depthwise-separable UNet with trilinear upsampling at three levels (L1/L2/L3). At each level we fuse $\widetilde{\mathbf{V}}^{\mathrm{Ca}}$ using a dual-path volumetric saliency and a Mixture-of-Experts (MoE) fuse; we term this module AMoE-3D, where A denotes attention implemented via the dual-path (channel & spatial) saliency gates.
Dual-path volumetric saliency (channel & spatial gates):
$\left\{\begin{aligned} \mathbf{A}_{c}\;&=\;\sigma\!\Big(\mathrm{MLP}(\mathrm{GAP}(\mathbf{X}))+\mathrm{MLP}(\mathrm{GMP}(\mathbf{X}))\Big),\\ \mathbf{A}_{s}\;&=\;\sigma\!\Big(g^{7\times 7\times 7}\!\left([\mathrm{Avg}(\mathbf{X});\,\mathrm{Max}(\mathbf{X})]\right)\Big),\\ \mathbf{Y}\;&=\;\mathbf{X}\odot\mathbf{A}_{c}\odot\mathbf{A}_{s}.\end{aligned}\right$. (12)
MoE-Fuse3D with gradient-energy gating:
$\left\{\begin{aligned} &\mathrm{GradEnergy3D}(\mathbf{Y})\;=\;\sum_{a\in\{x,y,z\}}\big\|\nabla_{a}\mathbf{Y}\big\|_{2}^{2},\\ &\boldsymbol{\alpha}\;=\;\operatorname{softmax}\!\big(W_{g}*\mathrm{GradEnergy3D}(\mathbf{Y})\big),\\ &\widetilde{\mathbf{Y}}\;=\;\mathbf{Y}+\sum_{k=1}^{K}\alpha_{k}\,E_{k}(\mathbf{Y}),\end{aligned}\right$. (13)
where $E_{k}$ are $1{\times}1{\times}1$ Conv–GELU–Conv experts. Here, $\mathbf{X}\!\in\!\mathbb{R}^{C\times D\times H\times W}$ is the level-$\ell$ Cartesian feature and $\widetilde{\mathbf{V}}^{\mathrm{Ca}}$ is the polar volume resampled onto the same Cartesian grid; $\mathbf{A}_{c}\!\in\!\mathbb{R}^{C\times 1\times 1\times 1}$ and $\mathbf{A}_{s}\!\in\!\mathbb{R}^{1\times D\times H\times W}$ are the channel and spatial gates. $\nabla_{a}$ denotes the discrete 3D gradient along axis $a\!\in\!\{x,y,z\}$, * is 3D convolution, and [;][\,;\,] concatenates along channels.
Why it helps.
Hierarchical polar injections aggregate far-field azimuthal cues at coarse levels—e.g., ring-consistent road/building/vegetation context —and refine near-field contact geometry at fine levels. Gradient-energy gating emphasizes high-contrast structures at category transitions (vehicles, poles) while suppressing overfitting on large ground flats (road), improving stability of foothold decisions.
Table 4: Ablations on QuadOcc (Panorama only). Core columns show if a component is enabled (✓) or disabled (✗).
Variant Core Designs Per-class IoU (QuadOcc) Overall
GDC Gait Disp. Compensation
DP–ER Dual Projection (Equi. + Raw)
BGV Bi–Grid Voxelization
AMoE–3D Attention–MoE 3D Fusion
$\blacksquare$ vehicle (0.90%)
$\blacksquare$ pedestrian (2.09%)
$\blacksquare$ road (52.34%)
$\blacksquare$ building (14.00%)
$\blacksquare$ vegetation (22.82%)
$\blacksquare$ terrain (7.85%)
SC Precision
SC Recall
SC IoU
SSC mIoU
(Q0) baseline 8.15 1.59 55.66 12.88 26.10 10.78 62.10 69.28 48.69 19.19
(Q1) + GDC 11.76 2.48 53.73 15.07 22.95 11.51 64.05 66.08 48.78 19.58
(Q2) + DP–ER 11.82 2.51 54.16 15.39 23.40 12.08 65.42 65.01 48.38 19.89
(Q3) + BGV 12.07 2.77 54.29 15.76 24.19 12.74 66.56 64.70 48.68 20.30
(Q4) + AMoE–3D 12.16 2.86 54.41 16.03 24.91 13.01 66.69 64.74 48.92 20.56
3.7 Segmentation Head and Loss
Method.
A $1{\times}1{\times}1$ head outputs per-voxel logits $\mathbf{Z}$. We supervise only valid voxels (from lifting visibility) with deep supervision at strides $\{1,2,4\}$.
Loss.
We follow MonoScene [cao2022monoscene] for most losses: (i) standard cross-entropy on valid voxels, (ii) the scene-class affinity (SCAL) terms for semantics and geometry, and (iii) the frustums proportion (FP) loss. We do not use the relation loss, as we observed it encourages co-occurrence priors that over-smooth azimuthal boundaries and suppress small near-field classes under panoramic legged setting, yielding no gain and occasional degradation on our data. The final objective is:
$\mathcal{L}_{\mathrm{total}}\;=\;\mathcal{L}_{\mathrm{CE}}\;+\;\mathcal{L}_{\mathrm{SCAL}}^{\mathrm{sem}}\;+\;\mathcal{L}_{\mathrm{SCAL}}^{\mathrm{geo}}\;+\;\mathcal{L}_{\mathrm{FP}}.$ (14)
Class re-weighting for $\mathcal{L}_{\mathrm{CE}}$ follows standard SSC practice.
4 Experiments
Table 5: Performance comparison on the QuadOcc validation set across different lighting conditions. Results are shown for day, dusk, and night scenarios. P (Precision), R (Recall), IoU and mIoU are reported. Methods marked with $\dagger$} use a sensor modality different from the original paper. The best performance for each input modality is highlighted in bold.
Method Day [Uncaptioned image] Dusk [Uncaptioned image] Night [Uncaptioned image]
P R IoU mIoU P R IoU mIoU P R IoU mIoU
LiDAR-based
SSCNet [song2017semantic] 67.12 62.95 48.12 14.03 52.97 76.34 45.50 14.88 57.81 75.30 48.60 9.51
SSCNet-full [song2017semantic] 80.56 58.52 51.28 16.04 73.79 64.34 52.37 17.14 73.64 67.31 54.24 11.56
LMSCNet [roldao2020lmscnet] 81.18 62.50 54.59 17.33 75.34 70.92 57.56 18.91 75.68 72.67 58.91 13.40
OccRWKV [wang2024occrwkv] 72.84 63.83 51.56 16.25 66.43 59.30 45.63 15.42 74.10 58.31 48.44 12.61
Vision-based
OccFormer [zhang2023occformer] 46.23 77.08 40.64 12.82 41.79 79.45 37.72 14.46 49.49 72.18 41.56 10.21
[…]
SGN [mei2023camera] 61.54 68.71 48.06 17.94 54.07 69.06 43.53 15.19 46.83 52.29 32.81 10.58
MonoScene [cao2022monoscene] 65.47 69.94 51.09 18.58 47.62 77.10 41.72 15.14 51.35 66.47 40.78 14.20
OneOcc (Ours) 69.10 64.66 50.15 21.15 57.46 71.98 46.96 19.86 63.56 53.75 41.09 13.50
4.1 Setup
Datasets. We evaluate on two panoramic occupancy benchmarks for legged/humanoid platforms. QuadOcc: real first-person $360^{\circ}$ data on a quadruped within a campus; day/dusk/night, sequence-heterogeneous split; 2424K consecutive frames with stride-5 training; $6$ classes on a $64{\times}64{\times}8$ grid (0.4 m). Human360Occ (H3O): CARLA-based human-ego $360^{\circ}$ with simulated gait over $16$ maps and diverse weather/lighting; RGB, depth, occupancy (two resolutions), and poses. We report within-city (H3O-Homo) and cross-city (H3O-Heter). See Table 2 and Table 3.
Metrics. Following semantic scene completion (SSC) practice, we report per-class IoU and mean IoU (mIoU) on non-empty voxels, plus overall precision (P) and recall (R).
Baselines. We re-train popular LiDAR SSC systems (SSCNet [song2017semantic], LMSCNet [roldao2020lmscnet], OccRWKV [wang2024occrwkv]) and adapt representative vision SSC methods (OccFormer [zhang2023occformer], VoxFormer [li2023voxformer], SGN [mei2023camera], MonoScene [cao2022monoscene]) to panoramas via calibrated unwrapping to equirectangular. For P+L entries in Table 2, we add LiDAR to the corresponding vision model per the table footnote.
Implementation. Unless noted, training follows official configs with minimal dataset-specific changes; schedules/image sizes/voxelization match for fairness. On H3O we disable the raw-projection branch (native equirectangular), keeping components as in ablations. Training uses a single NVIDIA RTX-4090 GPU.
Figure 4: Qualitative comparisons on (a) QuadOcc and (b) Human360Occ (val sets). Left to right: Reference panorama, ground-truth occupancy, baselines, and OneOcc (ours). Across both real (QuadOcc) and simulated (H3O) panoramic settings, OneOcc better preserves the global layout and $360^{\circ}$ continuity, produces cleaner class boundaries, and recovers occluded structures. Yellow dotted circles highlight typical improvements over LMSCNet/SGN-S and MonoScene (e.g., foreground-background disambiguation and far-range layout).
4.2 Quantitatively Comparisons
QuadOcc. Table 2 reports QuadOcc (val). OneOcc attains 20.56 mIoU, surpassing the best LiDAR baseline LMSCNet (18.44) and the strongest vision baseline MonoScene (19.19). SGN (P+L) achieves the highest geometry IoU (49.16) among vision entries by leveraging LiDAR, yet camera-only OneOcc delivers the best mIoU (20.56) among vision methods. Task-aligned panoramic fusion (DP-ER, BGV, AMoE-3D, with GDC on legged data) enables a camera-only pipeline to rival and even surpass popular LiDAR stacks at this range/resolution.
Human360Occ (H3O). Table 3 summarizes generalization: On H3O-Homo OneOcc reaches 37.29 mIoU (+3.83 vs. MonoScene), while on H3O-Heter it attains 32.23 mIoU (+8.08). The larger margin under heterogeneous cross-city shifts highlights OneOcc’s strong distribution robustness: DP-ER and AMoE-3D transfer across maps and lighting/weather, and their distortion-aware priors mitigate panoramic aliasing, yielding state-of-the-art camera-only mIoU under distribution shift.
Robustness to Lighting Conditions. Table 5 reports day/dusk/night. OneOcc leads the best vision baseline in day (21.15 vs. 18.58) and dusk (19.86 vs. 15.14); at night its mIoU (13.50) trails MonoScene (14.20) but shows higher precision, likely from frustum-artifact suppression (cf. Fig. 4). LiDAR trend. Dusk often exceeds day due to weaker solar background in near-IR, while night drops with altered return statistics and sparser echoes on low-albedo/specular surfaces, hurting long-range completion.
4.3 Ablations
We ablate GDC, DP-ER, BGV, AMoE-3D (Table 4). Starting from Q0 (ER-only, no GDC, single-grid, no AMoE-3D; 19.19 mIoU), we add: +GDC $\rightarrow$ 19.58 (+0.39), +DP-ER $\rightarrow$ 19.89 (+0.31), +BGV $\rightarrow$ 20.30 (+0.41), +AMoE-3D (full) $\rightarrow$ 20.56 (+0.26). Gains are additive: GDC stabilizes sampling; DP-ER provides complementary cues; BGV reduces discretization bias; AMoE-3D sharpens edges/contacts without over-smoothing flats.
4.4 Qualitative Comparisons
Fig. 4 shows typical cases. On QuadOcc, MonoScene exhibits frustum-direction artifacts (ghost elongation), while OneOcc suppresses them via DP-ER (cross-view corroboration), GDC (phase-stable sampling), and BGV (near/far balance). On Human360Occ, OneOcc recovers pedestrians close in color to background that competing vision methods miss. Overall, OneOcc preserves global layout and $360^{\circ}$ continuity and better handles occlusions.
4.5 Efficiency
Inference on an RTX 4090 at $608{\times}1216$, batch size $1$. OneOcc averages 69.93 ms (FPS $\approx$ 14.30). The model has 101.76M parameters and peaks at 1.86 GB CUDA allocation (2.65 GB reserved) in FP32. With mixed precision (FP32+FP16), latency drops to 52.84 ms (FPS $\approx$ 18.92), and peak allocation is 1.49 GB. The footprint remains moderate, making OneOcc suitable for onboard legged perception on a single GPU. Jetson Thor extrapolation: using published tensor-core budgets (RTX 4090 FP16 tensor 165 TFLOPS vs. Thor 517 TFLOPS at FP16 with 2:4 sparsity), a linear scale suggests >>10 FPS is achievable on Thor.
5 Conclusion
We introduced OneOcc, a panoramic camera-only framework for semantic occupancy on legged and humanoid platforms. OneOcc stabilizes feature phases under periodic motion via Gait Displacement Compensation (GDC), fuses raw annular and equirectangular streams with Dual-Projection (DP–ER), balances near–far geometry through Bi-Grid Voxelization (BGV), and aggregates multi-scale evidence using a hierarchical AMoE-3D decoder. Together, these components turn monocular panoramas into a geometry-aware 3D grid. We also contribute QuadOcc (real $360^{\circ}$ quadruped) and report cross-domain evaluation on Human360Occ, facilitating systematic study across robot morphologies and environments.
Limitations & Outlook. OneOcc assumes accurate calibration with bounded drift. A possible remedy is online extrinsic self-calibration, optionally regularized by lightweight odometry priors. We propose robotic occupancy as a intermediate layer for world models and vision-language-action models: tokens for language control, and occupancy-sequence pretraining for cross-robot transfer.

### v2

@@
 We also release two panoramic occupancy benchmarks: QuadOcc (real quadruped, first-person $360^{\circ}$) and Human360Occ (H3O) (CARLA human-ego $360^{\circ}$ with RGB, Depth, semantic occupancy; standardized within-/cross-city splits).
-OneOcc sets new state-of-the-art (SOTA): on QuadOcc it beats strong vision baselines and popular LiDAR ones; on H3O it gains +3.83 mIoU (within-city) and +8.08 (cross-city).
+OneOcc sets a new state of the art on QuadOcc, outperforming strong vision baselines and remaining competitive with classical LiDAR baselines; on H3O it gains +3.83 mIoU (within-city) and +8.08 (cross-city).
 Modules are lightweight, enabling deployable full-surround perception for legged/humanoid robots.
-Datasets and code will be publicly available at https://github.com/MasterHow/OneOcc.
+Datasets and code will be publicly available at OneOcc.
 1 Introduction
@@
 (a) Single-sensor panoramic semantic scene completion (SSC) on a quadruped: an omnidirectional camera observes full $360^{\circ}$ with gait-induced motion; OneOcc outputs voxelized semantic occupancy.
-(b) On QuadOcc, OneOcc reaches 20.56 mIoU, surpassing the best LiDAR baseline (LMSCNet, $18.44$) and vision baseline (MonoScene, $19.19$).
+(b) Summary on QuadOcc and Human360Occ (H3O): bars show mIoU(%).
+On QuadOcc, OneOcc reaches 20.56 mIoU, surpassing the best LiDAR baseline (LMSCNet, $18.44$) and vision baseline (MonoScene, $19.19$).
 On H3O, OneOcc attains 37.29/32.23 mIoU under within-/cross-city, outperforming the best vision baselines ($33.46/24.15$).
-A holistic 3D understanding of nearby surroundings is fundamental for robotic autonomy in navigation [wang2024omega, hoeller2022neural, wang2024he_nav, wang2024agrnav, patel2025hierarchical, li2024stereonavnet], locomotion [niijima2025realtime_multi_plane, dong2025marg, escontrela2025gaussgym, chen2025learning_traversal], manipulation [kartmann2021semantic, dengler2025efficient_manipulation, li2025manipdreamer3d, wang2025odyssey], and human-robot interaction [wu2025embodiedocc, wang2025embodiedocc++, liu2024revisit_human_scene_interaction].
-Among 3D perception tasks, Semantic Scene Completion (SSC)—predicting complete volumetric geometry and semantics from partial observations—has become central [lai2024rtonet, cao2024slcf_sequential, yan2024pointssc, zhang2025roboocc], with most progress driven by autonomous driving where forward-facing sensors on stable wheeled platforms yield benign motion and sensing.
+A holistic 3D understanding of nearby surroundings is fundamental for robotic autonomy in navigation [82, 28, 83, 84, 67, 39], locomotion [61, 19, 21, 11], manipulation [33, 18, 44, 86], and human-robot interaction [93, 81, 48].
+Among 3D perception tasks, Semantic Scene Completion (SSC)—predicting complete volumetric geometry and semantics from partial observations—has become central [36, 5, 99, 108], with most progress driven by autonomous driving where forward-facing sensors on stable wheeled platforms yield benign motion and sensing.
 Data Domain
 Datasets Camera Sur. Plat. Move. Env. Weath. Lighting Modality Volume Size Scenes Frames Anno.
-NYUv2 [nyu] Pinhole [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Indoor [Uncaptioned image] [Uncaptioned image] C&D $(144{\times}240{\times}240)$ 1.4K 1.4K [Uncaptioned image]
-ScanNet [scannet] Pinhole [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Indoor [Uncaptioned image] [Uncaptioned image] C&D $(31{\times}62{\times}62)$ 1.5K 1.5K [Uncaptioned image]
-SUNCG [song2017semantic] Pinhole [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Indoor [Uncaptioned image] [Uncaptioned image] C&D $(144{\times}240{\times}240)$ 46K 140K [Uncaptioned image]
-SemanticPOSS [semanticposs] n.a. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Outdoor [Uncaptioned image] [Uncaptioned image] L $(32{\times}256{\times}256)$ 6 3K [Uncaptioned image]
-SemanticKITTI [semantickitti] Pinhole [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Outdoor [Uncaptioned image] [Uncaptioned image] C&L $(32{\times}256{\times}256)$ 22 9K [Uncaptioned image]
-nuScenes [wang2023openoccupancy] Pinhole+Fisheye [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Outdoor [Uncaptioned image] [Uncaptioned image][Uncaptioned image][Uncaptioned image] C&L $(40{\times}512{\times}512)$ 850 200K [Uncaptioned image]
-Occ3D-Waymo [tian2024occ3d] Pinhole+Fisheye [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Outdoor [Uncaptioned image] [Uncaptioned image][Uncaptioned image][Uncaptioned image] C&L $(32{\times}200{\times}200)$ 1K 200K [Uncaptioned image]
-SSCBench-KITTI-360 [li2024sscbench] Fisheye [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Outdoor [Uncaptioned image] [Uncaptioned image] C&L $(32{\times}200{\times}200)$ 11 91K [Uncaptioned image]
-ScribbleSC [wang2024label_efficient] Pinhole [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Outdoor [Uncaptioned image] [Uncaptioned image] C&L $(32{\times}256{\times}256)$ 11 9K [Uncaptioned image]
-OmniHD-Scenes [zheng2024omnihd] Pinhole [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Outdoor [Uncaptioned image] [Uncaptioned image][Uncaptioned image] C&L&R 200 60K [Uncaptioned image][Uncaptioned image]
-WildOcc [zhai2024wildocc] Pinhole [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Outdoor [Uncaptioned image] [Uncaptioned image] C&L $(40{\times}100{\times}100)$ 5 10K [Uncaptioned image]
-DSEC-SSC [guo2025event] Pinhole [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Outdoor [Uncaptioned image] [Uncaptioned image][Uncaptioned image] C&L&E $(16{\times}128{\times}128)$ 6 3K [Uncaptioned image]
-ORAD-3D [min2025advancing_orad] Pinhole [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Outdoor [Uncaptioned image] [Uncaptioned image][Uncaptioned image][Uncaptioned image] C&L 145 58K [Uncaptioned image]
-Co3SOP [wu2025synthetic_v2x] Pinhole [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Outdoor [Uncaptioned image] [Uncaptioned image] C&L $(70{\times}1000{\times}1000)$ [Uncaptioned image]
+NYUv2 [75] Pinhole [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Indoor [Uncaptioned image] [Uncaptioned image] C&D $(144{\times}240{\times}240)$ 1.4K 1.4K [Uncaptioned image]
+ScanNet [16] Pinhole [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Indoor [Uncaptioned image] [Uncaptioned image] C&D $(31{\times}62{\times}62)$ 1.5K 1.5K [Uncaptioned image]
+SUNCG [76] Pinhole [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Indoor [Uncaptioned image] [Uncaptioned image] C&D $(144{\times}240{\times}240)$ 46K 140K [Uncaptioned image]
+SemanticPOSS [66] n.a. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Outdoor [Uncaptioned image] [Uncaptioned image] L $(32{\times}256{\times}256)$ 6 3K [Uncaptioned image]
+SemanticKITTI [2] Pinhole [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Outdoor [Uncaptioned image] [Uncaptioned image] C&L $(32{\times}256{\times}256)$ 22 9K [Uncaptioned image]
+nuScenes [89] Pinhole+Fisheye [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Outdoor [Uncaptioned image] [Uncaptioned image][Uncaptioned image][Uncaptioned image] C&L $(40{\times}512{\times}512)$ 850 200K [Uncaptioned image]
+Occ3D-Waymo [80] Pinhole+Fisheye [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Outdoor [Uncaptioned image] [Uncaptioned image][Uncaptioned image][Uncaptioned image] C&L $(32{\times}200{\times}200)$ 1K 200K [Uncaptioned image]
+SSCBench-KITTI-360 [42] Fisheye [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Outdoor [Uncaptioned image] [Uncaptioned image] C&L $(32{\times}200{\times}200)$ 11 91K [Uncaptioned image]
+ScribbleSC [88] Pinhole [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Outdoor [Uncaptioned image] [Uncaptioned image] C&L $(32{\times}256{\times}256)$ 11 9K [Uncaptioned image]
+OmniHD-Scenes [112] Pinhole [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Outdoor [Uncaptioned image] [Uncaptioned image][Uncaptioned image] C&L&R 200 60K [Uncaptioned image][Uncaptioned image]
+WildOcc [103] Pinhole [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Outdoor [Uncaptioned image] [Uncaptioned image] C&L $(40{\times}100{\times}100)$ 5 10K [Uncaptioned image]
+DSEC-SSC [26] Pinhole [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Outdoor [Uncaptioned image] [Uncaptioned image][Uncaptioned image] C&L&E $(16{\times}128{\times}128)$ 6 3K [Uncaptioned image]
+ORAD-3D [58] Pinhole [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Outdoor [Uncaptioned image] [Uncaptioned image][Uncaptioned image][Uncaptioned image] C&L 145 58K [Uncaptioned image]
+Co3SOP [92] Pinhole [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Outdoor [Uncaptioned image] [Uncaptioned image] C&L $(70{\times}1000{\times}1000)$ [Uncaptioned image]
 Human360Occ (ours) Panoramic [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Outdoor [Uncaptioned image] [Uncaptioned image][Uncaptioned image][Uncaptioned image] C&D $(16{\times}128{\times}128)$ 160 8K [Uncaptioned image]
@@
 Panoramic SSC with a single sensor.
-A single panoramic camera provides compact, true $360^{\circ}$ coverage but introduces annular distortions and seam artifacts after unwrapping [lin2025one_flight]—issues poorly handled by perspective-camera SSC pipelines.
+A single panoramic camera provides compact, true $360^{\circ}$ coverage but introduces annular distortions and seam artifacts after unwrapping [45]—issues poorly handled by perspective-camera SSC pipelines.
 This motivates a design that preserves raw panoramic cues while exploiting grid-friendly projections.
@@
 To advance panoramic SSC on legged/humanoid robots, we release two complementary benchmarks.
-QuadOcc is a real first-person $360^{\circ}$ dataset on a quadruped within a campus domain, with standardized day/dusk/night coverage and semi-automatic ground truth (multi-frame LiDAR aggregation, Grounded-SAM [ren2024grounded] initialization, targeted manual fixes).
+QuadOcc is a real first-person $360^{\circ}$ dataset on a quadruped within a campus domain, with standardized day/dusk/night coverage and semi-automatic ground truth (multi-frame LiDAR aggregation, Grounded-SAM [70] initialization, targeted manual fixes).
 It contains 10 scenes and 24K frames, uses 6 semantic categories for training.
@@
 It supports within-city (per-map 8:28{:}2 train/val) and cross-city (default 12/412{/}4 maps) splits.
+See details on the construction pipeline and dataset statistics in the supplementary material.
 Results in brief.
@@
 Legged autonomy motivates perception tailored to agile motion and strict payload budgets.
-Prior efforts include vision-based terrain reconstruction [zhang2022vision], large-scale datasets [patel2025tartanground], and traversability learning [deng2021vision, aegidius2025watch, oh2024trip]; complementary modules cover robust tracking [cao2025siamese, xin2024robotic] and instruction grounding [mei2024quadrupedgpt, han2025space].
-With increasing evidence that limited FoV hinders agility [li2024move, cheng2024quadruped, wang2025omni], panoramic sensors are adopted for omnidirectional tracking [luo2025omnidirectional], multimodal odometry [li2025limo], and even satellite communication on the move [liu2025starlink].
-Some works study probabilistic semantic mapping [chen2025particle], and humanoid systems explore panoramic/occupancy perception [cui2025humanoid, zhang2025humanoidpano, zhang2025roboocc, zhang2025occupancy_world_models_robots].
+Prior efforts include vision-based terrain reconstruction [104], large-scale datasets [68], and traversability learning [17, 1, 62]; complementary modules cover robust tracking [6, 95] and instruction grounding [57, 27].
+With increasing evidence that limited FoV hinders agility [41, 13, 90], panoramic sensors are adopted for omnidirectional tracking [52], multimodal odometry [40], and even satellite communication on the move [46].
+Some works study probabilistic semantic mapping [9], and humanoid systems explore panoramic/occupancy perception [15, 106, 108, 109].
 However, dense 3D semantic occupancy from a single panoramic camera on dynamic legged platforms is still largely open.
@@
 Method.
-We calibrate the omnidirectional panoramic annular lens (PAL) camera with the Taylor polynomial model [scaramuzza2006toolbox].
+We calibrate the omnidirectional panoramic annular lens (PAL) camera with the Taylor polynomial model [72].
 The unwrapping samples the raw annulus at coordinates given by the calibrated inverse mapping:
@@
 LiDAR-based
-SSCNet [song2017semantic] L 0.00 0.04 44.34 16.06 22.41 4.77 63.42 65.82 47.71 14.60
-SSCNet-full [song2017semantic] L 0.71 0.04 55.14 14.93 19.66 7.76 78.81 60.03 51.69 16.38
-LMSCNet [roldao2020lmscnet] L 0.88 0.20 57.02 16.45 24.39 11.70 79.70 64.49 55.40 18.44
-OccRWKV [wang2024occrwkv] L 2.05 0.35 54.17 10.29 24.92 8.38 71.98 62.80 50.47 16.69
+SSCNet [76] L 0.00 0.04 44.34 16.06 22.41 4.77 63.42 65.82 47.71 14.60
+SSCNet-full [76] L 0.71 0.04 55.14 14.93 19.66 7.76 78.81 60.03 51.69 16.38
+LMSCNet [71] L 0.88 0.20 57.02 16.45 24.39 11.70 79.70 64.49 55.40 18.44
+OccRWKV [85] L 2.05 0.35 54.17 10.29 24.92 8.38 71.98 62.80 50.47 16.69
 Vision-based
-VoxFormer-S [li2023voxformer] P+D 0.31 1.67 36.83 6.48 7.09 3.81 45.96 69.95 38.38 9.36
-SGN [mei2023camera] P+D 7.39 1.88 53.35 9.06 21.95 9.39 59.23 67.41 46.05 17.17
-VoxFormer-S [li2023voxformer] P+L 0.24 1.66 38.51 10.90 11.88 4.45 51.29 75.84 44.08 11.27
-VoxFormer-T [li2023voxformer] P+L 0.30 1.65 40.21 14.48 15.86 4.64 56.94 77.10 48.70 12.86
-SGN [mei2023camera] P+L 11.62 2.47 53.06 15.60 25.67 9.91 73.70 59.62 49.16 19.72
-OccFormer [zhang2023occformer] P 0.29 0.37 49.46 10.36 15.00 2.64 45.79 76.99 40.28 13.02
-MonoScene [cao2022monoscene] P 8.15 1.59 55.66 12.88 26.10 10.78 62.10 69.28 48.69 19.19
+VoxFormer-S [43] P+D 0.31 1.67 36.83 6.48 7.09 3.81 45.96 69.95 38.38 9.36
+SGN [56] P+D 7.39 1.88 53.35 9.06 21.95 9.39 59.23 67.41 46.05 17.17
+VoxFormer-S [43] P+L 0.24 1.66 38.51 10.90 11.88 4.45 51.29 75.84 44.08 11.27
+VoxFormer-T [43] P+L 0.30 1.65 40.21 14.48 15.86 4.64 56.94 77.10 48.70 12.86
+SGN [56] P+L 11.62 2.47 53.06 15.60 25.67 9.91 73.70 59.62 49.16 19.72
+OccFormer [107] P 0.29 0.37 49.46 10.36 15.00 2.64 45.79 76.99 40.28 13.02
+MonoScene [4] P 8.15 1.59 55.66 12.88 26.10 10.78 62.10 69.28 48.69 19.19
 OneOcc (Ours) P 12.16 2.86 54.41 16.03 24.91 13.01 66.69 64.74 48.92 20.56
@@
 Precision Recall IoU mIoU
-VoxFormer-[li2023voxformer] P+Dpred. 22.27 17.44 2.86 5.65 0.87 0.14 0.00 0.00 61.46 0.18 40.21 59.49 31.57 11.09
-VoxFormer-[li2023voxformer] P+Dgt 35.84 25.86 12.32 9.90 3.73 0.70 0.19 0.17 64.79 0.93 59.91 71.12 48.19 15.44
-SGN-[mei2023camera] P+Dpred. 54.55 44.61 15.51 35.16 23.91 7.68 4.52 15.77 65.24 19.46 61.82 64.36 46.05 28.64
-SGN-[mei2023camera] P+Dpred. 54.81 47.44 15.64 33.22 24.23 8.53 2.62 13.51 65.88 19.72 62.28 63.72 45.98 28.56
-OccFormer [zhang2023occformer] P 54.11 46.41 15.14 32.84 17.57 4.30 0.00 1.65 65.08 11.72 57.30 69.94 45.98 24.88
-MonoScene [cao2022monoscene] P 62.45 57.18 19.97 38.91 29.76 13.77 5.03 16.76 70.15 20.67 68.78 71.66 54.08 33.46
+VoxFormer-[43] P+Dpred. 22.27 17.44 2.86 5.65 0.87 0.14 0.00 0.00 61.46 0.18 40.21 59.49 31.57 11.09
+VoxFormer-[43] P+Dgt 35.84 25.86 12.32 9.90 3.73 0.70 0.19 0.17 64.79 0.93 59.91 71.12 48.19 15.44
+SGN-[56] P+Dpred. 54.55 44.61 15.51 35.16 23.91 7.68 4.52 15.77 65.24 19.46 61.82 64.36 46.05 28.64
+SGN-[56] P+Dpred. 54.81 47.44 15.64 33.22 24.23 8.53 2.62 13.51 65.88 19.72 62.28 63.72 45.98 28.56
+OccFormer [107] P 54.11 46.41 15.14 32.84 17.57 4.30 0.00 1.65 65.08 11.72 57.30 69.94 45.98 24.88
+MonoScene [4] P 62.45 57.18 19.97 38.91 29.76 13.77 5.03 16.76 70.15 20.67 68.78 71.66 54.08 33.46
 OneOcc (Ours) P 63.82 62.92 22.86 41.03 33.74 17.79 6.47 26.58 71.13 26.55 69.14 74.18 55.73 37.29
@@
 Precision Recall IoU mIoU
-VoxFormer-[li2023voxformer] P+Dpred. 19.02 15.82 2.84 5.86 0.95 0.10 0.00 0.01 61.62 0.10 36.68 52.00 27.40 10.63
-VoxFormer-[li2023voxformer] P+Dgt 32.55 25.95 8.98 9.93 4.68 0.44 0.00 0.00 65.45 0.26 59.08 61.21 42.99 14.82
-SGN-[mei2023camera] P+Dpred. 42.57 23.94 9.42 15.84 18.22 10.21 1.55 3.11 63.44 11.93 45.84 47.11 30.26 20.02
-SGN-[mei2023camera] P+Dpred. 43.47 24.70 9.19 20.27 17.21 8.21 1.85 5.08 61.10 14.89 47.82 50.98 32.76 20.60
-OccFormer [zhang2023occformer] P 47.00 39.01 10.14 24.85 11.84 1.52 0.05 0.77 66.60 6.93 53.12 63.57 40.73 20.87
-MonoScene [cao2022monoscene] P 49.68 41.41 11.04 20.57 20.34 10.36 1.98 2.79 69.68 13.69 67.39 55.00 43.44 24.15
+VoxFormer-[43] P+Dpred. 19.02 15.82 2.84 5.86 0.95 0.10 0.00 0.01 61.62 0.10 36.68 52.00 27.40 10.63
+VoxFormer-[43] P+Dgt 32.55 25.95 8.98 9.93 4.68 0.44 0.00 0.00 65.45 0.26 59.08 61.21 42.99 14.82
+SGN-[56] P+Dpred. 42.57 23.94 9.42 15.84 18.22 10.21 1.55 3.11 63.44 11.93 45.84 47.11 30.26 20.02
+SGN-[56] P+Dpred. 43.47 24.70 9.19 20.27 17.21 8.21 1.85 5.08 61.10 14.89 47.82 50.98 32.76 20.60
+OccFormer [107] P 47.00 39.01 10.14 24.85 11.84 1.52 0.05 0.77 66.60 6.93 53.12 63.57 40.73 20.87
+MonoScene [4] P 49.68 41.41 11.04 20.57 20.34 10.36 1.98 2.79 69.68 13.69 67.39 55.00 43.44 24.15
 OneOcc (Ours) P 58.60 48.28 16.34 25.98 30.68 19.35 12.87 14.16 72.89 23.11 67.89 64.77 49.58 32.23
@@
 $\Delta_{s}\;=\;\mathrm{Linear}_{0}\!\Big(\mathrm{GAP}(\mathbf{F}^{v}_{1/s})\Big),$ (9)
-where $\mathrm{Linear}_{0}$ has all weights and bias set to zero so the initial warp is identity (cf. zero-conv [zhang2023adding]). Before lifting we correct and resample:
+where $\mathrm{Linear}_{0}$ has all weights and bias set to zero so the initial warp is identity (cf. zero-conv [105]). Before lifting we correct and resample:
 $\widehat{\mathbf{p}}\;$ $=\;\pi_{v}(\mathbf{c};\boldsymbol{\kappa})\;+\;\Delta_{s},$ (10)
 […]
+In practice, this also upgrades the lifting operator from integer-index gather [4] to bilinear sampling, which reduces projection aliasing and further stabilizes GDC.
 Here, $v\!\in\!\{\text{unwrapped},\text{spherical}\}$ indexes the projection path; $\pi_{v}(\mathbf{c};\boldsymbol{\kappa})$ projects a 3D voxel centroid $\mathbf{c}$ to image pixels using the calibrated Taylor camera with parameters $\boldsymbol{\kappa}$.
@@
 Why it helps.
-GDC routes the phase error back to 2D, preventing mis-attributed voxels. The zero-init head stabilizes early training and leaves features unperturbed when motion is tiny ($\Delta_{s}\!\approx\!0$).
+GDC routes the phase error back to 2D, preventing mis-attributed voxels.
+The zero-init head stabilizes early training and leaves features unperturbed when motion is tiny ($\Delta_{s}\!\approx\!0$).
+We additionally provide qualitative analysis of GDC in the supplementary material.
 3.6 Hierarchical AMoE-3D
@@
 Loss.
-We follow MonoScene [cao2022monoscene] for most losses: (i) standard cross-entropy on valid voxels, (ii) the scene-class affinity (SCAL) terms for semantics and geometry, and (iii) the frustums proportion (FP) loss.
+We follow MonoScene [4] for most losses: (i) standard cross-entropy on valid voxels, (ii) the scene-class affinity (SCAL) terms for semantics and geometry, and (iii) the frustums proportion (FP) loss.
 We do not use the relation loss, as we observed it encourages co-occurrence priors that over-smooth azimuthal boundaries and suppress small near-field classes under panoramic legged setting, yielding no gain and occasional degradation on our data.
@@
 LiDAR-based
-SSCNet [song2017semantic] 67.12 62.95 48.12 14.03 52.97 76.34 45.50 14.88 57.81 75.30 48.60 9.51
-SSCNet-full [song2017semantic] 80.56 58.52 51.28 16.04 73.79 64.34 52.37 17.14 73.64 67.31 54.24 11.56
-LMSCNet [roldao2020lmscnet] 81.18 62.50 54.59 17.33 75.34 70.92 57.56 18.91 75.68 72.67 58.91 13.40
-OccRWKV [wang2024occrwkv] 72.84 63.83 51.56 16.25 66.43 59.30 45.63 15.42 74.10 58.31 48.44 12.61
+SSCNet [76] 67.12 62.95 48.12 14.03 52.97 76.34 45.50 14.88 57.81 75.30 48.60 9.51
+SSCNet-full [76] 80.56 58.52 51.28 16.04 73.79 64.34 52.37 17.14 73.64 67.31 54.24 11.56
+LMSCNet [71] 81.18 62.50 54.59 17.33 75.34 70.92 57.56 18.91 75.68 72.67 58.91 13.40
+OccRWKV [85] 72.84 63.83 51.56 16.25 66.43 59.30 45.63 15.42 74.10 58.31 48.44 12.61
 Vision-based
-OccFormer [zhang2023occformer] 46.23 77.08 40.64 12.82 41.79 79.45 37.72 14.46 49.49 72.18 41.56 10.21
+OccFormer [107] 46.23 77.08 40.64 12.82 41.79 79.45 37.72 14.46 49.49 72.18 41.56 10.21
 […]
 […]
-SGN [mei2023camera] 61.54 68.71 48.06 17.94 54.07 69.06 43.53 15.19 46.83 52.29 32.81 10.58
-MonoScene [cao2022monoscene] 65.47 69.94 51.09 18.58 47.62 77.10 41.72 15.14 51.35 66.47 40.78 14.20
+SGN [56] 61.54 68.71 48.06 17.94 54.07 69.06 43.53 15.19 46.83 52.29 32.81 10.58
+MonoScene [4] 65.47 69.94 51.09 18.58 47.62 77.10 41.72 15.14 51.35 66.47 40.78 14.20
 OneOcc (Ours) 69.10 64.66 50.15 21.15 57.46 71.98 46.96 19.86 63.56 53.75 41.09 13.50
@@
 We report within-city (H3O-Homo) and cross-city (H3O-Heter).
-See Table 2 and Table 3.
+See supplementary for the data collection pipelines and statistics of QuadOcc and H3O.
 Metrics. Following semantic scene completion (SSC) practice, we report per-class IoU and mean IoU (mIoU) on non-empty voxels, plus overall precision (P) and recall (R).
 Baselines.
-We re-train popular LiDAR SSC systems (SSCNet [song2017semantic], LMSCNet [roldao2020lmscnet], OccRWKV [wang2024occrwkv]) and adapt representative vision SSC methods (OccFormer [zhang2023occformer], VoxFormer [li2023voxformer], SGN [mei2023camera], MonoScene [cao2022monoscene]) to panoramas via calibrated unwrapping to equirectangular.
+We re-train popular LiDAR SSC systems (SSCNet [76], LMSCNet [71], OccRWKV [85]) and adapt representative vision SSC methods (OccFormer [107], VoxFormer [43], SGN [56], MonoScene [4]) to panoramas via calibrated unwrapping to equirectangular.
 For P+L entries in Table 2, we add LiDAR to the corresponding vision model per the table footnote.
+Details in the supplementary.
 Implementation.
@@
 On H3O we disable the raw-projection branch (native equirectangular), keeping components as in ablations.
-Training uses a single NVIDIA RTX-4090 GPU.
+Training uses a single RTX-4090 GPU.
+Details in the supplementary.
 Figure 4: Qualitative comparisons on (a) QuadOcc and (b) Human360Occ (val sets).
@@
 Gains are additive: GDC stabilizes sampling; DP-ER provides complementary cues; BGV reduces discretization bias; AMoE-3D sharpens edges/contacts without over-smoothing flats.
+We further provide more detailed per-module ablations, calibration-robustness analysis, temporal aggregation baselines, and resolution scaling results in the supplementary material.
 4.4 Qualitative Comparisons
@@
 OneOcc averages 69.93 ms (FPS $\approx$ 14.30).
-The model has 101.76M parameters and peaks at 1.86 GB CUDA allocation (2.65 GB reserved) in FP32.
+The model has 101.76M parameters and peaks at 1.82 GB CUDA allocation (2.58 GB reserved) in FP32.
 With mixed precision (FP32+FP16), latency drops to 52.84 ms (FPS $\approx$ 18.92), and peak allocation is 1.49 GB.
 The footprint remains moderate, making OneOcc suitable for onboard legged perception on a single GPU.
-Jetson Thor extrapolation: using published tensor-core budgets (RTX 4090 FP16 tensor 165 TFLOPS vs.
-Thor 517 TFLOPS at FP16 with 2:4 sparsity), a linear scale suggests >>10 FPS is achievable on Thor.
+We additionally report measured on-device runtime and a module-wise latency breakdown on NVIDIA Jetson AGX Orin in the supplementary material.
 5 Conclusion
@@
 A possible remedy is online extrinsic self-calibration, optionally regularized by lightweight odometry priors.
-We propose robotic occupancy as a intermediate layer for world models and vision-language-action models: tokens for language control, and occupancy-sequence pretraining for cross-robot transfer.
+We propose robotic occupancy as an intermediate layer for world models and vision-language-action models: tokens for language control, and occupancy-sequence pretraining for cross-robot transfer.
+Acknowledgment
+This research was funded by the Natural Science Foundation of Zhejiang Province (Grant No.
+LZ24F050003), the National Natural Science Foundation of China (Grant Nos.
+12174341 and 62473139), the Hunan Provincial Research and Development Project (Grant No.
+2025QK3019), the MOE AcRF Tier 1 SSHR-TG Incubator Grant FY24 (Grant No.
+RSTG7/24), and the opening project of the State Key Laboratory of Autonomous Intelligent Unmanned Systems (Grant No.
+ZZKF2025-2-10).
+Hao Shi was also supported by the China Scholarship Council (CSC) for joint Ph.D.
+training at Nanyang Technological University (NTU).